Tell what the output value is for the function machine for the given values.
(a) 2
(b) 10
(c) -4
^2 + 1
1
Expert's answer
2013-04-08T08:07:05-0400
Tell what the output value is for the function machine for the given values:
f(x)=x^2 + 1
&
(a) x=2
f(x)=x^2 + 1
We need to substitute x = 2:
f(x=2) = 2^2+1=4+1=5
Answer: 5
&
(b) x=10
f(x)=x^2 + 1
We need to substitute x = 10:
f(x=10)=10^2+1=100+1=101
Answer: 101
&
(c) x=-4
f(x)=x^2 + 1
We need to substitute x = -4:
f(x=-4)=(-4)^2+1=16+1=17
Answer: 17